OpenStudy (anonymous):

Multiply. (3x – 5)2 9x2 – 30x + 25 6x2 – 10 9x2 – 25 9x2 – 15x + 25

OpenStudy (anonymous):

(the twos are exponents fyi)

OpenStudy (compassionate):

9x^2 -25

OpenStudy (compassionate):

Multiple by the 2nd square of everything in the brackets.

OpenStudy (ajprincess):

\((a-b)^2=a^2-2ab+b^2\) Here a=3x and b=5

OpenStudy (anonymous):

thank you!!!!

OpenStudy (ajprincess):

yw:)